the following costs were incurred by odd a mfg cost

DM........105200

DL.........221400

MOH.....98600

based on the above data

a.determine prime cost

b.conversion cost

c.mfg cost


1
Expert's answer
2021-08-20T06:39:36-0400

Solution:

a.). Prime cost = Direct Materials + Direct Labor

= 105,200 + 221,400 = 326,600

 

b.). Conversion cost = Direct labor + Manufacturing Overheads (MOH)

= 221,400 + 98,600 = 320,000

 

c.). Manufacturing (mfg) cost = Direct Materials + Direct Labor + Manufacturing Overheads

= 105,200 + 221,400 + 98,600 = 425,200